OpenStudy (anonymous):

Show all work to simplify (2)/(x)+(2)/(x+1)-(2)/(x+2) . Use complete sentences to explain how to simplify this expression. Remember to list all restrictions.

OpenStudy (callisto):

(2)/(x)+(2)/(x+1)-(2)/(x+2) = (2)(x+1)(x+2) / x(x+1)(x+2) + 2(x)(x+2)/x(x+1)(x+2) - 2x(x+1)/x(x+1)(x+2) = (2x^2 +6x +4) / x(x+1)(x+2) +(2x^2 +4x) /x(x+1)(x+2) - (2x^2 +2x)/x(x+1)(x+1) = [(2x^2 +6x +4) +(2x^2 +4x) - (2x^2 +2x)]/x(x+1)(x+2) = (2x^2 +8x +4)/x(x+1)(x+2) = 2(x^2 +4x+2) /x(x+1)(x+2) restriction neither x nor x+1 nor x+2 =0

OpenStudy (anonymous):

thank you . (:

OpenStudy (callisto):

be careful and check my steps, i'd easily make mistakes when dealing with too many terms..

OpenStudy (anonymous):

okay i will , im so far behind in school right now and i just need the most help i can take . so i really appreciate it . (: